# Repetita Iuvant: Data Repetition Allows SGD to Learn High-Dimensional Multi-Index Functions

### v1

@@ A.1 Proof of Theorem 2
There exists a constant $c(\delta)$ such that if $\gamma_{0}(\delta)\leq c(\delta)$, $m_{0}>\kappa(\delta)/\sqrt{d}$ and
$T\leq c(\delta)\gamma^{-2}d,$
we have
$\mathbb{P}\left(m_{t}\geq\frac{m_{0}}{2}+\frac{1}{2}\sum_{s=0}^{t}\mathbb{E}[\langle\Phi(\bm{z}_{s}),\bm{w}^{\star}\rangle]\quad\forall t\leq T\right)\geq 1-\frac{\delta}{2}$
It remains to study the above expectation. By the computations done in the proof sketch, we have
$\langle\Phi(\bm{z}_{t}),\bm{w}^{\star}\rangle=h^{\star}(\langle\bm{w}^{\star},\bm{z}_{t}\rangle)\sigma^{\prime}\left(\langle\bm{w}_{t},\bm{z}_{t}\rangle+\rho h^{\star}(\langle\bm{w}^{\star},\bm{z}_{t}\rangle)\sigma^{\prime}(\langle\bm{w}_{t},\bm{z}_{t}\rangle)\cdot\left\lVert\bm{z}_{t}\right\rVert^{2}\right)\langle\bm{w^{\star}},\bm{z}_{t}\rangle.$
We show that this is well approximated by $\phi(m_{t})$:
Lemma 3.
Deterministically, the following bound holds:
$\left|\mathbb{E}\left[\langle\Phi(\bm{z}_{t}),\bm{w}^{\star}\rangle\right]-\phi(m)\right|\leq\frac{C}{\sqrt{d}}$ (18)
@@
The distribution of the pair $(\langle\bm{w}_{t},\bm{z}_{t}\rangle,\langle\bm{w^{\star}},\bm{z}_{t}\rangle)$ is the same as the one of $(x,x^{\star})$, so the only difference between both expressions is the replacement of $\left\lVert\bm{z}_{t}\right\rVert^{2}$ by $d$. Since $\sigma^{\prime\prime}$ is bounded, we can write
$\mathbb{E}\left[\left|\langle\Phi(\bm{z}_{t}),\bm{w}^{\star}\rangle-\phi(m)\right|\right]$ $\leq C\rho\mathbb{E}\left[\left|h^{\star}(\langle\bm{w}^{\star},\bm{z}_{t}\rangle)^{2}\sigma^{\prime}(\langle\bm{w}_{t},\bm{z}_{t}\rangle)\left(\left\lVert\bm{z}_{t}\right\rVert^{2}-d\right)\langle\bm{w^{\star}},\bm{z}_{t}\rangle\right|\right]$
$\leq C^{\prime}\rho\sqrt{\mathbb{E}\left[\left(\left\lVert\bm{z}_{t}\right\rVert^{2}-d\right)^{2}\right]}$
$\leq\frac{C^{\prime\prime}}{\sqrt{d}},$
where at the last line we used that $\rho=O(d^{-1})$. ∎
Finally, we can obtain the desired difference equation:
@@
There exists a constant $c(\delta)$ such that if $\gamma_{0}(\delta)\leq c(\delta)$, $m_{0}>\kappa(\delta)/\sqrt{d}$ and
$T\leq c(\delta)\gamma^{-2}d^{-1}=:\bar{T},$
we have for large enough $d$
$\mathbb{P}\left(m_{t}\geq\frac{m_{0}}{4}+\frac{1}{2}\sum_{s=0}^{t}\phi(m)\quad\forall t\leq T\right)\geq 1-\frac{\delta}{2}$
Proof.
From Lemmas 2 and 3,
$\mathbb{P}\left(m_{t}\geq\frac{m_{0}}{2}+\frac{\gamma}{2}\sum_{s=0}^{t}\phi(m)-\frac{C\gamma T}{\sqrt{d}}\quad\forall t\leq T\right)\geq 1-\frac{\delta}{2}.$
The lemma is therefore proven as soon as
$\frac{C\gamma T}{\sqrt{d}}\leq\frac{m_{0}}{4},\quad\text{or}\quad T\leq C^{\prime}\gamma^{-1}\sqrt{d}m_{0}.$
But since $\sqrt{d}m_{0}\geq\kappa(\delta)$, and $\gamma\ll 1$, this condition follows from $T\leq c(\delta)\gamma^{-2}d$ when $d$ is large enough. ∎
Proving Theorem 2 will therefore ensue from the study of the function $\phi$. In particular, the results of Ben Arous et al. [2021] imply the following theorem:
Theorem 4.
There exists a choice of $\eta>0$ and $\epsilon$ in Assumption 3, depending on $\phi$, such that the following holds:
  • If $\phi(0)\neq 0$, then letting $\gamma=\gamma_{0}(\delta)d^{-1}$, with probability $1-\delta$,
    $t_{\eta}^{+}\leq C(\delta)d,$
  • If $\phi(0)=0$ and $\phi^{\prime}(0)\neq 0$, then letting $\gamma=\gamma_{0}(\delta)(d\log(d))^{-1}$, with probability $1-\delta$,
    $t_{\eta}^{+}\leq C(\delta)d\log(d)^{2}$
Proof.
We first treat the case where $\phi(0)\neq 0$. Upon changing the choice of $\epsilon$, we can assume that $\phi(0)$ has the same sign as $m_{0}$, and is therefore positive without loss of generality. Let $c=\phi(0)/2$; we define $\eta$ to be a small constant such that $\phi(m)>c$ on $[0,\eta]$. Then, by Proposition 1, we have
$\mathbb{P}\left(m_{0}\geq\frac{\kappa(\delta)}{\sqrt{d}},\quad m_{t}\geq\frac{m_{0}}{4}+\frac{c\gamma}{2}t\quad\forall t\leq\bar{T}\wedge t_{\eta}^{+}\right)\geq 1-\delta$ (19)
As a result, $t_{\eta}^{+}\leq 2\eta/{c\gamma}$ as long as the latter bound is lower than $\bar{T}$. The above condition is equivalent to
$\gamma_{0}(\delta)\leq\frac{c(\delta)\cdot c\eta}{2},$
which can be ensured by decreasing $\gamma_{0}(\delta)$.
The case where $\phi(0)$ is treated similarly: by the discrete Grönwall inequality,
$\mathbb{P}\left(m_{0}\geq\frac{\kappa(\delta)}{\sqrt{d}},\quad m_{t}\geq\frac{m_{0}}{4}e^{c\gamma t}\quad\forall t\leq\bar{T}\wedge t_{\eta}^{+}\right)\geq 1-\delta,$ (20)
and thus $t_{\eta}^{+}\leq C(\delta)\gamma^{-1}\log(d)$, again under the condition that this bound is lower than $\bar{T}$. This time, the bound is equivalent to
$\gamma_{0}(\delta)\leq\frac{c(\delta)}{C(\delta)},$
which can again be easily ensured. ∎
Analysis of $\phi$
To prove Theorem 2, it remains to show the following proposition:
Proposition 2.
If $\ell_{p}^{\star}=1$, then for almost every choice of $\rho_{0}$, we have $\phi(0)\neq 0$. On the other hand, if $\ell_{p}^{\star}=2$, then $\phi^{\prime}(0)\neq 0$ for almost every choice of $\rho_{0}$.
The main ingredient here is to write $\phi(m)=\phi(m;\rho_{0})$ and differentiate w.r.t $\rho_{0}$. We show the following:
@@
$\left.\frac{\partial^{k}\phi(m;\rho_{0})}{\partial\rho_{0}^{k}}\right|_{\rho_{0}=0}=\mathbb{E}\left[h^{\star}(x^{\star})^{(k+1)}x^{\star}\sigma^{k+1}(x)\sigma^{\prime}(x)^{k}\right]$ (21)
@@
$\frac{\partial\phi}{\partial\rho_{0}}=\mathbb{E}\left[\sigma^{\prime}(x)h^{\star}(x^{\star})^{2}\sigma^{\prime\prime}(x+\rho_{0}\sigma^{\prime}(x)h^{\star}(x^{\star}))x^{\star}\right];$
@@
Then, by Proposition 11.37 from O’Donnell [2014], we have
$\left.\frac{\partial^{k}\phi(m;\rho_{0})}{\partial\rho_{0}^{k}}\right|_{\rho_{0}=0}=u_{0}^{(k)}v_{0}^{(k)}+u_{1}^{(k)}v_{1}^{(k)}\cdot m+o(m)$
@@
A.2 Proof of Lemma 1

### v2

@@ B.1 Proof of Theorem 2
The function $\Psi$ is approximately one-dimensional: there exists a function $\phi$ such that for any $\mathbf{w}\in\mathbb{S}^{d-1}$, we have
$|\langle\mathbf{w}^{\star},\Psi(\mathbf{w})\rangle-a_{0}\phi(\mathbf{w},\mathbf{w}^{\star})|\leq\frac{C}{d}$
@@
Define the following quantities:
$\Xi(x,x^{\star},x^{\bot})=h^{\star}(x^{\star})\sigma^{\prime}(x+a_{0}\rho_{0}\sigma^{\prime}(x)h^{\star}(x^{\star}))x^{\bot}$ (39)
$\phi(m)=\mathbb{E}[\Xi(x,x^{\star},x^{\bot})]\ \text{for}\ \dbinom{x}{x^{\star}}\sim\mathcal{N}\left(\mathbf{0},\begin{pmatrix}1&m\\ m&1\end{pmatrix}\right)\ \text{and}\ x^{\bot}=x^{\star}-mx.$ (40)
For simplicity, we let $u=\langle\mathbf{w},\mathbf{z}\rangle+a_{0}\rho_{0}h^{\star}(\langle\mathbf{w}^{\star},\mathbf{z}\rangle)$. Doing a Taylor expansion of $\sigma^{\prime}$ in the second order, we have
$\sigma^{\prime}\left(\langle\mathbf{w},\mathbf{z}\rangle+\rho h^{\star}(\langle\mathbf{w}^{\star},\mathbf{z}\rangle)\sigma^{\prime}(\langle\mathbf{w},\mathbf{z}\rangle)\cdot\left\lVert\mathbf{z}\right\rVert^{2}\right)=$ $\sigma^{\prime}(u)+\sigma^{\prime\prime}(u)\rho h^{\star}(\langle\mathbf{w}^{\star},\mathbf{z}\rangle)\sigma^{\prime}(\langle\mathbf{w},\mathbf{z}\rangle)\cdot\left(\left\lVert\mathbf{z}\right\rVert^{2}-d\right)$
$+R(u)\rho^{2}h^{\star}(\langle\mathbf{w}^{\star},\mathbf{z}\rangle)^{2}\sigma^{\prime}(\langle\mathbf{w},\mathbf{z}\rangle)^{2}\cdot\left(\left\lVert\mathbf{z}\right\rVert^{2}-d\right)^{2},$
where $R(u)$ is the Taylor remainder of $\sigma^{\prime}$ at second order around $u$.
The distribution of the triplet $(\langle\mathbf{w},\mathbf{z}\rangle,\langle\mathbf{w^{\star}},\mathbf{z}\rangle,\langle\mathbf{w^{\star}},\mathbf{z}^{\bot}\rangle)$ is the same as the one of $(x,x^{\star},x^{\bot})$, hence
$\mathbb{E}\left[\langle\Phi(\mathbf{z}),\mathbf{w}^{\star}\rangle\right]-a_{0}\phi(m)=$ $\underbrace{\rho\mathbb{E}\left[\sigma^{\prime\prime}(u)h^{\star}(\langle\mathbf{w}^{\star},\mathbf{z}\rangle)^{2}\sigma^{\prime}(\langle\mathbf{w},\mathbf{z}\rangle)\cdot\left(\left\lVert\mathbf{z}\right\rVert^{2}-d\right)\langle\mathbf{w}^{\star},\mathbf{z}^{\bot}\rangle\right]}_{(1)}$
$+\underbrace{\rho^{2}\mathbb{E}\left[R(u)h^{\star}(\langle\mathbf{w}^{\star},\mathbf{z}\rangle)^{3}\sigma^{\prime}(\langle\mathbf{w},\mathbf{z}\rangle)^{2}\cdot\left(\left\lVert\mathbf{z}\right\rVert^{2}-d\right)^{2}\langle\mathbf{w}^{\star},\mathbf{z}^{\bot}\rangle\right]}_{(2)}.$
We first bound (2). From the Hölder inequality (since $\frac{1}{2}+\frac{1}{8}+\frac{1}{8}+\frac{1}{8}+\frac{1}{8}=1$),
$(2)$ $\leq\rho^{2}\mathbb{E}\left[R(u)^{8}\right]^{\frac{1}{8}}\mathbb{E}\left[h^{\star}(x^{\star})^{8}\right]^{\frac{1}{8}}\mathbb{E}\left[\sigma^{\prime}(x)^{8}\right]^{\frac{1}{8}}\mathbb{E}\left[\langle\mathbf{w}^{\star},\mathbf{z}^{\bot}\rangle^{8}\right]^{\frac{1}{8}}\mathbb{E}\left[\left(\left\lVert\mathbf{z}\right\rVert^{2}-d\right)^{4}\right]^{\frac{1}{2}}$
Since $\sigma$ is three times differentiable almost everywhere, and from the polynomial bounds of Assumption 4, the first four expectations are bounded by an absolute constant. The last term is the centered fourth moment of a $\chi^{2}(d)$ distribution, which is equal to $12d(d+4)$. As a result, using that $\rho^{2}=O(d^{-2})$,
$(2)\leq Cd^{-1}.$
For the first term, we let $\mathbf{z}=\mathbf{z}_{\parallel}+\mathbf{z}^{\prime}$, where $\mathbf{z}_{\parallel}$ is the projection of $\mathbf{z}$ on the span of $\mathbf{w}$ and $\mathbf{w}^{\star}$. We can then write
$(1)$ $=\underbrace{\mathbb{E}\left[\sigma^{\prime\prime}(u)\rho h^{\star}(\langle\mathbf{w}^{\star},\mathbf{z}\rangle)^{2}\sigma^{\prime}(\langle\mathbf{w},\mathbf{z}\rangle)\cdot\left\lVert\mathbf{z}_{\parallel}\right\rVert^{2}\cdot\langle\mathbf{w}^{\star},\mathbf{z}^{\bot}\rangle\right]}_{(1^{\prime})}$
$\quad+\underbrace{\mathbb{E}\left[\sigma^{\prime\prime}(u)\rho h^{\star}(\langle\mathbf{w}^{\star},\mathbf{z}\rangle)^{2}\sigma^{\prime}(\langle\mathbf{w},\mathbf{z}\rangle)\cdot\left(\left\lVert\mathbf{z^{\prime}}\right\rVert^{2}-d\right)\langle\mathbf{w}^{\star},\mathbf{z}^{\bot}\rangle\right]}_{(1^{\prime\prime})}.$
Up to the factor of $\rho$, the term $(1^{\prime})$ is the expectation of a polynomially bounded function of a $2$-dimensional Gaussian variable, hence
$(1^{\prime})\leq C\rho\leq C^{\prime}d^{-1}.$
In $(1^{\prime\prime})$, the vector $\mathbf{z^{\prime}}$ is independent from $\langle\mathbf{w},\mathbf{z}\rangle$ and $\langle\mathbf{w}^{\star},\mathbf{z}\rangle$, so the expectation factorizes as
$|(1^{\prime\prime})|$ $=\left|\mathbb{E}\left[\sigma^{\prime\prime}(u)\rho h^{\star}(\langle\mathbf{w}^{\star},\mathbf{z}\rangle)^{2}\sigma^{\prime}(\langle\mathbf{w},\mathbf{z}\rangle)\langle\mathbf{w}^{\star},\mathbf{z}^{\bot}\rangle\right]\cdot\mathbb{E}\left[\left(\left\lVert\mathbf{z^{\prime}}\right\rVert^{2}-d\right)\right]\right|$
$\leq C\rho|d-2-d|\leq C^{\prime}d^{-1},$
where we used that $\mathbf{z}^{\prime}$ is a $(d-2)$-dimensional Gaussian variable. This concludes the proof. ∎
Lemma 2 implies that Assumption 8 holds for $\Psi$ and any choice of $q$, as long as $d$ is large enough. As a result, the proof of Theorem 2 hinges on the following facts about $\phi$:
@@
For almost any choice of $\rho_{0}$ (w.r.t to the Lebesgue measure),
  • if $\ell_{p}^{\star}=1$, then $\phi(0)\neq 0$;
  • if $\ell_{p}^{\star}=2$, then $\phi^{\prime}(0)\neq 0$.
of Theorem 2.
Our goal is to apply Proposition 4 to the function $\phi$. If $\ell_{p}^{\star}=1$, then $\phi(0)\neq 0$ by the previous lemma, and with a uniform initialization of $\mathbf{w}^{\star}$ we have $m_{0}\phi(0)>0$ with probability $1/2$. Combining this with the $(1-\delta)$ probability event of Proposition 4 for small enough $\epsilon$, with probability $1-\delta$, we have $t_{\eta}\leq C(\delta)d^{-1}$ with probability $1/2-\delta$. We now move to the proof of the case where $\ell_{p}^{\star}=1$. In this case, Proposition 4 implies the desired result as soon as $\phi^{\prime}(0)>0$ (which corresponds in the $k=1$ setting to the only eigenvalue of $d\phi(0)$ being positive). Since the Lebesgue measure is invariant w.r.t multiplication by ±$\pm 1$, the factor $a_{0}\rho_{0}$ in the definition of $\phi^{\prime}(0)$ has the same distribution as $\rho_{0}$, hence we can write $\phi^{\prime}(0)=a^{\prime}_{0}|\phi^{\prime}(0)|$ where $a^{\prime}_{0}$ is also uniform in $\{-1,1\}$. The event $\phi^{\prime}(0)>0$ therefore also has probability $1/2$, and we conclude as before. ∎
Analysis of $\phi$: proof of Proposition 5
The main ingredient here is to write $\phi(m)=\phi(m;\rho_{0})$ and differentiate w.r.t $\rho_{0}$. For simplicity, we take $a_{0}=1$ in the definition of $\phi$. We show the following:
@@
$\left.\frac{\partial^{k}\phi(m;\rho_{0})}{\partial\rho_{0}^{k}}\right|_{\rho_{0}=0}=\mathbb{E}\left[h^{\star}(x^{\star})^{k+1}x^{\bot}\sigma^{(k+1)}(x)\sigma^{\prime}(x)^{k}\right]$ (41)
@@
$\frac{\partial\phi}{\partial\rho_{0}}=\mathbb{E}\left[\sigma^{\prime}(x)h^{\star}(x^{\star})^{2}\sigma^{\prime\prime}(x+\rho_{0}\sigma^{\prime}(x)h^{\star}(x^{\star}))x^{\bot}\right];$
@@
$\Delta^{(k)}$ $=\mathbb{E}\left[h^{\star}(x^{\star})^{k}\right]$
Then, by Proposition 11.37 from O’Donnell [2014], and using $x^{\bot}=x^{\star}-mx$:
$\left.\frac{\partial^{k}\phi(m;\rho_{0})}{\partial\rho_{0}^{k}}\right|_{\rho_{0}=0}$ $=\mathbb{E}\left[h^{\star}(x^{\star})^{k+1}x^{\star}\sigma^{(k+1)}(x)\sigma^{\prime}(x)^{k}\right]-m\cdot\mathbb{E}\left[h^{\star}(x^{\star})^{k+1}x\sigma^{(k+1)}(x)\sigma^{\prime}(x)^{k}\right]$
$=\left(u_{0}^{(k)}v_{0}^{(k)}+u_{1}^{(k)}(v_{1}^{(k)}+\Delta^{(k)})\cdot m+o(m)\right)-m\left(u_{1}^{(k)}\Delta^{(k)}+o(1)\right)$
$=u_{0}^{(k)}v_{0}^{(k)}+u_{1}^{(k)}v_{1}^{(k)}\cdot m+o(m)$
@@
B.2 Proof of Theorem 4
The proof of Theorem 4 follows in the same way as the one of Theorem 2. Recall the definition of $\Phi$ in eq. (38), and define
$\Xi(z,\mathbf{z}^{\star},\mathbf{z}^{\bot})=h^{\star}(\mathbf{z}^{\star})\sigma^{\prime}(z+a_{0}\rho_{0}\sigma^{\prime}(z)h^{\star}(\mathbf{z}^{\star}))\mathbf{z}^{\bot}$ (42)
$\psi(\mathbf{p})=\mathbb{E}\left[\Xi(z,\mathbf{z}^{\star},\mathbf{z}^{\bot})\right]\ \text{for}\ \dbinom{z}{\mathbf{z}^{\star}}\sim\mathcal{N}\left(\mathbf{0},\begin{pmatrix}1&\mathbf{p}\\ \mathbf{p}&I_{k}\end{pmatrix}\right)\ \text{and}\ \mathbf{z}^{\bot}=\mathbf{z}^{\star}-z\mathbf{p}.$ (43)
Again, the variables $(z,\mathbf{z}^{\star})$ correspond to $\langle\mathbf{w},\mathbf{z}\rangle,W^{\star}\mathbf{z}$, and the vector $\mathbf{p}$ to $W^{\star}\mathbf{w}$. As before, the only difference between $\mathbb{E}\left[\Phi(\mathbf{w})\right]$ and $\psi(W^{\star}\mathbf{w})$ lies in replacing $\left\lVert\mathbf{z}\right\rVert$ with $d$, and the proof of Lemma 2 implies the following: for any $\mathbf{w}\in\mathbb{S}^{d-1}$,
$\left\lVert W^{\star}\Psi(\mathbf{w})-\psi(W^{\star}\mathbf{w})\right\rVert\leq\frac{C}{d}.$
As a result, the function $\Psi$ is approximately $k$-dimensional for any choice of $q$ as $d\to\infty$.
We can view the vector $\psi(\mathbf{p})$ as a random variable depending on the choice of $a_{0},\rho_{0}$. Since $a_{0}$ is a sign, we will only consider the dependency of $\psi$ in $\rho_{0}$. As before, let $u_{i}^{(k)}$ be the $i$-th Hermite coefficient of $z\mapsto\sigma^{(k+1)}(z)\sigma^{\prime}(z)^{k}$, and $C_{i}^{(k)}$ the $i$-th Hermite tensor of $h^{\star}$ [Grad, 1949] – in particular, $C_{i}^{(k)}$ is a tensor of order $i$. We have the analogous result to Lemma 3:
Lemma 4.
Under Assumption 4, the function $\psi(\mathbf{p};\rho_{0})$ is analytic in $\rho_{0}$, with
$\left.\frac{\partial^{k}\psi(\mathbf{p};\rho_{0})}{\partial\rho_{0}^{k}}\right|_{\rho_{0}=0}=u_{0}^{(k)}C_{1}^{(k)}+u_{1}^{(k)}C_{2}^{(k)}\mathbf{p}+o(\left\lVert\mathbf{p}\right\rVert).$
Proof.
As in Lemma 3, we can write $\psi(\mathbf{0})=\psi(\mathbf{0},\rho_{0})$, with
$\left.\frac{\partial^{k}\psi(\mathbf{p};\rho_{0})}{\partial\rho_{0}^{k}}\right|_{\rho_{0}=0}=\mathbb{E}\left[h^{\star}(\mathbf{z}^{\star})^{k+1}\mathbf{z}^{\bot}\sigma^{(k+1)}(z)\sigma^{\prime}(z)^{k}\right].$
From [Dandi et al., 2024a, Lemma 19], we have
$\mathbb{E}\left[h^{\star}(\mathbf{z}^{\star})^{k+1}\mathbf{z}^{\star}\sigma^{(k+1)}(z)\sigma^{\prime}(z)^{k}\right]=\sum_{i=0}^{\infty}u_{i+1}^{(k)}\langle\mathbf{p}^{\otimes i},C_{i}^{(k)}\rangle\mathbf{p}+\sum_{i=0}^{\infty}u_{i}^{(k)}C_{i+1}^{(k)}\times_{1,\dots,i}\mathbf{p}^{\otimes i}.$
Since $\mathbf{z}^{\bot}=\mathbf{z}-z\mathbf{p}$, we find
$\left.\frac{\partial^{k}\psi(\mathbf{p};\rho_{0})}{\partial\rho_{0}^{k}}\right|_{\rho_{0}=0}=\sum_{i=0}^{\infty}u_{i}^{(k)}\left(C_{i+1}^{(k)}\times_{1,\dots,i}\mathbf{p}^{\otimes i}-\langle C_{i+1}^{(k)},\mathbf{p}^{\otimes(i+1)}\rangle\mathbf{p}\right).$ (44)
For $i\geq 0$, we have $\left\lVert C_{i+1}^{(k)}\times_{1,\dots,i}\mathbf{p}^{\otimes i}\right\rVert=O\left(\left\lVert\mathbf{p}\right\rVert^{i}\right)$ and $\langle C_{i}^{(k)},\mathbf{p}^{\otimes(i)}\rangle=O\left(\left\lVert\mathbf{p}\right\rVert^{i}\right)$. The only terms in the sum which are not of order $\left\lVert\mathbf{p}\right\rVert^{2}$ are the first two terms in the left sum, which yields the statement of the Lemma. ∎
We are now ready to show the first part of Theorem 4. We begin with a characterization of $U^{\star}$ whenever $\ell_{p}^{\star}=1$:
Lemma 5.
If $\ell_{p}^{\star}=1$, then
$U^{\star}=\operatorname{span}\left(\left\{C_{1}^{(k)}\right\}_{k\geq 1}\right).$
If $\ell_{p}^{\star}=2$, then
$U^{\star}=\bigoplus_{k\geq 1}\mathrm{Im}\left(C_{2}^{(k)}\right).$
Proof.
By the properties of the Hermite tensors, we have for any $v\in V^{\star}$ and any polynomial $p$
$\mathbb{E}\left[y^{k}H_{1}(\langle\mathbf{v},\mathbf{z}\rangle)\right]=\langle\mathbf{v},C_{1}^{(k)}\rangle$
As a result, if $\mathbf{v}=\sum_{k\geq 0}v_{k}C_{1}^{(k)}$, then by letting $p(y)=\sum_{k\geq 0}a_{k}y^{k}$ we have
$\mathbb{E}\left[p(y)H_{1}(\langle\mathbf{v},\mathbf{z}\rangle)\right]=\left\lVert\mathbf{v}\right\rVert^{2}\neq 0.$
Conversely, if $\mathbf{v}$ is orthogonal to every vector $C_{1}^{(k)}$, then $\mathbb{E}\left[y^{k}H_{1}(\langle\mathbf{v},\mathbf{z}\rangle)\right]=0$ for all $k\geq 0$, which implies that $\mathbf{v}\notin U^{\star}$.
The case $\ell_{p}^{\star}=2$ is similar, since this time
$\mathbb{E}\left[y^{k}H_{2}(\langle\mathbf{v},\mathbf{z}\rangle)\right]=\mathbf{v}^{\top}C_{2}^{(k)}\mathbf{v}.$
As a result, for fixed $k$, the largest vector space $U_{k}$ such that $\mathbb{E}\left[y^{k}H_{2}(\langle\mathbf{v},\mathbf{z}\rangle)\right]$ for all $\mathbf{v}\in U_{k}$ is $\ker(C_{2}^{(k)})=\mathrm{Im}(C_{2}^{(k)})^{\bot}$. The result ensues from $U^{\star}=\bigcap_{k\geq 1}U_{k}^{\bot}$. ∎
Using Assumption 4, the following corollary ensues:
Corollary 1.
If $\ell_{p}^{\star}=1$, then for any $\mathbf{v}\in U^{\star}$ and almost every $\rho_{0}$,
$\langle\mathbf{v},\psi(\mathbf{0})\rangle\neq 0.$
As a result, with probability one, the set $\{\psi(\mathbf{0};\rho_{0,i})\}_{i\in[p]}$ spans $U^{\star}$ as soon as $p\geq k$.
If $\ell_{p}^{\star}=2$, then with probability at least $1/2$ on the choice of $\rho_{0},a_{0}$, $d\psi(\mathbf{0})$ is symmetric and has a strictly positive eigenvalue.
Proof.
Let $\mathbf{v}\in U^{\star}$, and $k\geq 1$ such that $\langle\mathbf{v},C_{1}^{(k)}\rangle\neq 0$. Then
$\left.\frac{\partial^{k}\langle\mathbf{v},\psi(\mathbf{0};\rho_{0})\rangle}{\partial\rho_{0}^{k}}\right|_{\rho_{0}=0}=u_{0}^{(k)}\langle\mathbf{v},C_{1}^{(k)}\rangle\neq 0,$
so the function $\rho_{0}\mapsto\langle\mathbf{v},\psi(\mathbf{0};\rho_{0})\rangle$ is not the zero function. Since it is also analytic, it is nonzero for almost every choice of $\rho$.
We now prove that if $p\leq\dim(U^{\star})$, then the family $\{\psi(\mathbf{0};\rho_{0,i})\}_{i\in[p]}$ is free, which implies our second result. Assume by recursion that $\{\psi(\mathbf{0};\rho_{0,i})\}_{i\in[p-1]}$ is free, and let $\mathbf{v}_{p}\in U^{\star}$ orthogonal to this family. Then with probability one, $\langle\mathbf{v}_{p},\psi(\mathbf{0};\rho_{0,p})\rangle\neq 0$, so $\psi(\mathbf{0};\rho_{0,p})$ is independent from $\{\psi(\mathbf{0};\rho_{0,i})\}_{i\in[p-1]}$.
Finally, if $\ell_{p}^{\star}=2$, then by symmetry of the distribution of $\rho_{0}$ we can write $\psi(\mathbf{p};a_{0},\rho_{0})=a_{0}\psi(\mathbf{p};1,\rho^{\prime})$ with $\rho^{\prime}=a_{0}\rho_{0}$ chosen independently from $a_{0}$. Writing
$d\psi(\mathbf{0};\rho^{\prime})=\sum_{k\geq 1}(\rho^{\prime})^{k}u_{1}^{(k)}C_{2}^{(k)},$
we can see that $d\psi(\mathbf{0})$ is a symmetric matrix, which is non-zero with probability one on $\rho^{\prime}$. Since either $d\psi(\mathbf{0})$ or $-d\psi(\mathbf{0})$ has a positive eigenvalue, $a_{0}d\psi(\mathbf{0})$ does with probability $1/2$. ∎
We are now ready to prove Theorem 4.
of Theorem 4.
We begin with the case $\ell_{p}^{\star}=1$. Fix a $p=p(\delta)$ to be chosen later, and let $\psi_{i}(\mathbf{p})=\psi(\mathbf{p};a_{0,i},\rho_{0,i})$ for $i\in[p]$. With probability $1/2$ on $\mathbf{w}_{0,i}$, we have $\langle\mathbf{w}_{0,i},\psi_{i}(\mathbf{0})\rangle$; let $p_{+}$ be the number of such neurons. For $p=O(\log(1/\delta))$, we have $p_{+}>\max(p/4,k)$ with probability at least $1-\frac{\delta}{2}$.
In particular, with probability one, the set $\{\psi_{i}(\mathbf{0})\}_{i\in[p_{+}]}$ spans $U^{\star}$. Let $\epsilon>0$ to be chosen later, $\eta_{i}(\epsilon)$ adapted to $\psi_{i}$ in Proposition 4, $\tau_{i}$ the hitting time for $\eta_{i}$ in the $i$-th neuron. Consider the time $\tau=\min_{i}\tau_{i}$; applying Proposition 4 to $\delta^{\prime}=\delta/p$, we have
$\tau\leq\max_{i\in[p_{+}]}\frac{\eta_{i}(1+\sqrt{k}\epsilon)}{\gamma_{0}\left\lVert\psi_{i}(\mathbf{0})\right\rVert}d\leq C(\delta)d.$
However, we also have $\tau>c(\delta)\gamma^{-1}$, and hence from Proposition 2, for any $i\in[p_{+}]$
$\left\lVert\mathbf{w}_{t,i}\right\rVert\geq\gamma\tau\left\lVert\psi_{i}(\mathbf{0})\right\rVert-\sqrt{k}\eta_{i}\epsilon\geq c^{\prime}(\delta)$
for small enough $\epsilon$. This proves the first part of the theorem. Now, consider the vectors
$\mathbf{u}_{i}=\frac{\Pi^{\star}\mathbf{w}_{\tau,i}}{\left\lVert\Pi^{\star}\mathbf{w}_{\tau,i}\right\rVert};$
by the same argument as in Proposition 4, for any $i\in[p_{+}]$,
$\left\lVert\mathbf{u}_{i}-\frac{\psi_{i}(\mathbf{0})}{\left\lVert\psi_{i}(\mathbf{0})\right\rVert}\right\rVert\leq\sqrt{k}\epsilon.$
Since the $\{\psi_{i}(\mathbf{0})\}_{i\in[p_{+}]}$ span $U^{\star}$ with probability one, so do the $\{\mathbf{u}_{i}\}_{i\in[p_{+}]}$ for a small enough choice of $\epsilon$.
For the case $\ell_{p}^{\star}=2$, we define the $p_{+}$ “good” neurons as those with $d\psi_{i}(\mathbf{0})$ having a positive eigenvalue. Since this happens with probability at least $1/2$ independently for each neuron, the bounds on $p(\delta)$ from earlier still apply. For those neurons, and small enough $\epsilon$, we define the $\eta_{i}(\epsilon)$ as those in Proposition 4, and $\tau=\max_{i}\tau_{i}$. Then by Proposition 4, we have $\tau\leq C(\delta)d\log(d)^{2}$ for some constant $C(\delta)$.
Now, fix a vector $i\in[p_{+}]$. We define the $j$-th “excursion” to be the $j$-th sequence of times $\tau_{i}^{(j)},\dots,\tau_{i}^{(j)}+e_{i}^{(j)}$ such that $\left\lVert\Pi^{\star}\mathbf{w}_{t,i}\right\rVert\geq\eta_{i}$. Then:
  • With probability at least $1-e^{-c\gamma}$, we have $\left\lVert\mathbf{w}_{t+1,i}-\mathbf{w}_{t,i}\right\rVert\leq C\sqrt{\gamma}$, and thus for any excursion $j$ we have $\left\lVert\Pi^{\star}\mathbf{w}_{\tau_{i}^{(j)}+e_{i}^{(j)}+1,i}\right\rVert\geq\eta_{i}/2$.
  • between two excursions, the results of Proposition 3 apply, and thus the norm of $\left\lVert\Pi^{\star}\mathbf{w}_{t,i}\right\rVert$ increases exponentially.
As a result, for $\tau_{i}\leq t\leq\tau$, we have $\left\lVert\Pi^{\star}\mathbf{w}_{t,i}\right\rVert>\eta_{i}/2$, which finishes the proof of Theorem 4. ∎
Appendix C Additional proofs
C.1 Proof of Lemma 1